An inventor received a royalty payment from a licensee for $15,000.00 that he wants to invest over the next 35 years on an account that he expects to earn 7% per year. Calculate the total accumulation in the account at the end of this period, (a) using both the standard notation, and (b) the factor formula, and (c) checking with Excel.
Answer #1

(a) Future Value = P(1 + i)n = 15,000(1 + 0.07)35 = $160,149

(b) Future Value = P(F/P, i, n) = 15,000(F/P, 7%, 35) = 15,000(10.6766) = $160,149

(c)

Present value -15000
Interest rate 7%
Term(Year) 35
Compounding period per year 1
Future value = FV(7%,35,0,-15000, 1)

FV = FV(7%,35,0,-15000, 1) = $160,149